# AdaSemiCD: An Adaptive Semi-Supervised Change Detection Method Based on Pseudo-Label Evaluation

### v1

@@ IV-B2 Qualitative results
To provide a clearer illustration of the change detection capability of our proposed method, we conducted visualizations on three different datasets. Fig. 4 showcases an example of the visualizations on the WHU-CD test sets. It is apparent that on both datasets, our approach has notably mitigated the common issues of missed and false detections. In challenging scenarios, our method could still effectively identify the areas of change that were of interest to us. Regarding the image pairs in the lower part of Fig. 4, there is significant interference information present between the images. Many of the detected changes are unrelated to the buildings of interest. The absence of adequate supervision information makes it challenging to mitigate such interference, leading to decreased model performance. Additionally, the task is further complicated by the detection of small and densely changing areas, which proves to be difficult for the model.
Alternative semi-supervised models may have been trained to detect errors because of guidance issues in this complicated scenario. Consequently, it can be seen that in such instances the semi-supervised method performs even worse than the supervised method trained solely on labeled data. Our model, on the other hand, adaptively excludes these noises during the training phase and integrates parameters from the model that has completed its evolution through the training process. In these intricate cases, the quality of pseudo-labels improves progressively until achieving a reliable level, furnishing an accurate signal for model training. Therefore, it is evident that our model excels in these complex regions (highlighted by boxes in Fig. 4). This marks the foundation of our proposed method.
Figure 4: Visualizations of different models on the WHU-CD dataset. Pre and Post represent the image of the front and back temporal respectively, and GT is the Ground Truth.
@@ IV-C Ablation study
In this section, we conduct some ablation studies, mainly to verify the role of each module in our proposed approach and to explore some of our hyperparameter selections in the experiment.
@@
Hyperparameters in Ramp-up: The Ramp-up process has a significant influence on the performance of our AdaSemiCD on SSCD. Therefore, we conduct experiments on the selection of two hyperparameters ($\gamma$ and $w_{max}$) a and b that control the Ramp-up process. As shown in Table V, our method achieves the best performance on the three datasets under the combination of parameters (0.1, 10), (0.1, 0.1), and (0.1, 1.0) respectively. Moreover, our method is sensitive to this hyperparameter, and inappropriate parameter selection will cause large performance attenuation. This is because our method conducts supervised training on labeled samples and unsupervised training on unlabeled samples at the same time. If the relationship between the two cannot be properly balanced, overfitting on labeled samples or excessive noise interference from unlabeled samples will be caused. All of our remaining experiments were performed at this hyperparameter setting, and the hyperparameters of the compared methods were consistent with the best choices in their original paper.
@@ IV-D Complexity Analysis
Since we utilize the same architecture with FPA and RCR, the number of training parameters (46.85M) and computational amount (585.85 GFLOPs) were the same. The variations in parameters, FLOPs, and the time required for training and inference are presented in Table VI. Our AdaSemiCD is comparable to other methods with reference count, FLOPs, and inference time. The main reason for the longer training time was that it took about 0.3s for each iteration to generate and evaluate pseudo-labels twice, while it only took about 0.006s and 0.03s for fusion and EMA parameters updating, respectively. Moreover, our model balances performance and time consumption, offering a notable performance benefit with only a slight increase in time.
@@
IV-E Limitation Analysis and Future Work
Although our approach is effective in most scenarios, we also found in our experiments that there are certain limitations in some cases. As shown in Fig. 5, our approach struggled and even performed worse than the supervised approach when faced with the detection of areas of leptosomatic variation, such as highways, rivers, and so on, especially from the example in the upper right corner, it can be seen that our method has great advantages in the identification and detection of blocky change areas. This is because when our AdaFusion transforms samples, it is the fusion of patch to patch, which easily truncates this changing region that throughout the entire image, and there is too little real data with annotations, resulting in no effective knowledge to be learned.
Therefore, in future research work, more targeted fusion methods could designed to avoid such information loss. For example, a changing region is regarded as an instance and sample fusion is carried out at the instance level. In this way, the characteristics of different change types can be preserved to a great extent while noise is removed, additional supervisory information is introduced and sample diversity is expanded. We will also actively explore the application of this adaptive mechanism to other semi-supervised tasks and frameworks.
Figure 5: Examples of common error cases in the CDD dataset.
@@ V Conclusion
In this study, we present AdaSemiCD, a flexible semi-supervised framework for change detection. This framework assesses the quality of pseudo-labels on unlabeled training samples and implements adaptive modifications based on the assessment outcomes, which include sample fusion (AdaFusion), and parameter updates (AdaEMA). Despite the complexity of the scenes, our model successfully identifies the areas of interest with minimal interference during training. Empirical evidence from several publicly available CD datasets attests to the efficacy of our methodology. Looking ahead, this adaptive processing technique holds promise for potential application in other semi-supervised tasks.

### v2

@@ IV-B2 Qualitative Results
Fig. 4 and Fig. 5 showcase some examples of the visualizations on the test sets of building and mutilclass CD datasets respectively, in which the area selected in the box is the error-prone area. It is apparent that on those datasets, our approach has notably mitigated the common issues of missed and false detections. In challenging scenarios, our method could still effectively identify the areas of change that were of interest to us.
Some failure cases of the detected changes are unrelated to the buildings of interest. The absence of adequate supervision information makes it challenging to mitigate such interference, leading to decreased model performance. Additionally, the task is further complicated by the detection of small and densely changing areas, which proves to be difficult for the model.
[…]
In complex scenarios, alternative semi-supervised models may encounter guidance issues, leading them to inadvertently amplify errors during training. As a result, semi-supervised methods may perform worse than supervised methods trained exclusively on labeled data in such situations. Our model, however, adaptively mitigates these noises during the training phase by dynamically excluding them and integrating parameters from a progressively refined model throughout the training process. In these intricate cases, the quality of pseudo-labels is incrementally improved until they reach a reliable standard, thereby providing an accurate signal for model training. Consequently, our model demonstrates superior performance in these challenging regions, as highlighted by the boxes in Fig. 4 and Fig. 5. This adaptive approach forms the cornerstone of our proposed method.
@@ IV-C Ablation Study
Pseudo-label qualification metric: As described in Section III-B, the pseudo-label evaluation metric we proposed is based on information entropy, enhanced by class rebalancing and confusion region amplification. To demonstrate the effectiveness, we first evaluated our AdaSemiCD using information entropy as the sole metric, then incrementally incorporated class rebalancing and confusion region amplification, and finally evaluated the full model, which integrates all components.
The experimental results are presented in Table V. When only information entropy is used as the evaluation metric for implementing the adaptive training mechanism, significant performance improvements are achieved, attributed to the AdaFusion and AdaEMA modules we designed. Incorporating class rebalancing and confusion region amplification based on information entropy further enhances performance, demonstrating that these improvements enable more accurate identification of pseudo-labels during training, thereby facilitating more precise adaptive operations. Notably, the gain from class rebalancing exceeds that from confusion region amplification. This is because, with class rebalancing, the evaluation of pseudo labels places greater emphasis on the foreground, somewhat neglecting background identification, thus allowing AdaFusion to focus more on the uncertainty region in the foreground. In addition, confusion region amplification proves more beneficial as more labeled data becomes available. Finally, using uncertainty, an overall assessment that includes both components, leads to further performance gains, suggesting that these two improvements are independent and can be effectively combined for more accurate assessments.
TABLE V: Ablation study of pseudo-label qualification metric on LEVIR-CD Dataset.
Method 5% 10% 20% 40%
$IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA
Sup. only 61.0 97.60 66.8 98.13 72.3 98.44 74.9 98.60
Entropy 73.2 (+12.2) 98.50 (+0.90) 77.4 (+10.6) 98.75 (+0.62) 78.6 (+6.3) 98.81 (+0.37) 79.2 (+4.3) 98.89 (+0.29)
Entropy+rebalance 76.0 (+15.0) 98.63 (+1.03) 78.7 (+11.9) 98.82 (+0.69) 79.6 (+7.3) 98.85 (+0.41) 79.6 (+4.7) 98.86 (+0.26)
Entropy+confusion 74.6 (+13.6) 98.61 (+1.01) 78.0 (+11.2) 98.79 (+0.66) 79.2 (+6.9) 98.82 (+0.38) 79.8 (+4.9) 98.87 (+0.27)
Uncertainty 77.7 (+16.7) 98.78 (+1.18) 79.4 (+12.6) 98.87 (+0.74) 80.3 (+8.0) 98.92 (+0.48) 80.6 (+5.7) 98.93 (+0.33)
We have stored the pseudo-labels generated during the training process, and Fig. 6 illustrates the $IoU^{c}$ between the pseudo-labels and the corresponding Ground Truth throughout training, both with and without class rebalancing in the pseudo-label evaluation (note: the Ground Truth used here for unlabeled samples are solely for calculating the $IoU^{c}$ with the pseudo-labels and are not involved in any other aspect of the training process). It is evident that, although class rebalancing does not directly improve the pseudo-labels, it enables a more accurate evaluation of the pseudo-labels. As a result, our adaptive training mechanism enhances the quality of them, bringing them closer to the real labels, which leads to a higher $IoU^{c}$ between them.
Figure 6: The influence of class rebalancing on the quality of pseudo-labels generated for unlabeled samples during training.
@@ IV-D Complexity Analysis
Since we utilize the same architecture with FPA and RCR, the number of training parameters (46.85M) and computational amount (585.85 GFLOPs) were the same. The variations in parameters, FLOPs, and the time required for training and inference are presented in Table VII. Our AdaSemiCD is comparable to other methods with reference count, FLOPs, and inference time. The main reason for the longer training time was that it took about 0.3s for each iteration to generate and evaluate pseudo-labels twice, while it only took about 0.006s and 0.03s for fusion and EMA parameters updating, respectively. Moreover, our model balances performance and time consumption, offering a notable performance benefit with only a slight increase in time.
@@
Hyperparameters in Ramp-up: The Ramp-up process has a significant influence on the performance of our AdaSemiCD on SSCD. Therefore, we conduct experiments on the selection of two hyperparameters ($\gamma$ and $w_{max}$) that control the Ramp-up process. As shown in Table VI, our method achieves the best performance on the three datasets under the combination of parameters (0.1, 10), (0.1, 0.1), and (0.1, 1.0) respectively. Moreover, our method is sensitive to this hyperparameter, and inappropriate parameter selection will cause large performance attenuation. This is because our method conducts supervised training on labeled samples and unsupervised training on unlabeled samples at the same time. If the relationship between the two cannot be properly balanced, overfitting on labeled samples or excessive noise interference from unlabeled samples will be caused. All of our remaining experiments were performed at this hyperparameter setting, and the hyperparameters of the compared methods were consistent with the best choices in their original paper.
@@ V Conclusion
In this study, we present AdaSemiCD, a flexible semi-supervised framework for change detection. This framework assesses the quality of pseudo-labels on unlabeled training samples and implements adaptive modifications based on the assessment outcomes, which include sample fusion (AdaFusion), and parameter updates (AdaEMA). Despite the complexity of the scenes, our model successfully identifies the areas of interest with minimal interference during training. Empirical evidence from ten publicly available CD datasets attests to the efficacy of our methodology. Looking ahead, this adaptive processing technique holds promise for potential application in other semi-supervised tasks.